OpenStudy (anonymous):

A circle has its center at (-2, -3) and a radius of 3 units. What is the equation of the circle? (x + 2)2 + (y - 3)2 = 9 (x - 2)2 + (y - 3)2 = 3 (x + 2)2 + (y + 3)2 = 9 (x + 2)2 + (y + 3)2 = 3

OpenStudy (anonymous):

@robtobey please help

OpenStudy (acxbox22):

what do you think is the answer?

OpenStudy (cwrw238):

the general form of the equation is (x - a)^2 + (y - b)^2 = r^2 where center = (a,b) and radius = r

OpenStudy (anonymous):

I dont understand this

OpenStudy (anonymous):

its either a or c

OpenStudy (cwrw238):

so if you compare this with the above options you'll find the correct one

OpenStudy (anonymous):

is it c?

OpenStudy (cwrw238):

well substitute a = -2 and b = -3 and r = 3 into the general form and you'll get the correct option

OpenStudy (cwrw238):

lets try the first part (x - a)^2 a = -2 , so we have (x -(-2))^2 = (x + 2)^2 now do the same for b and r

OpenStudy (cwrw238):

do you see what i am doing?

OpenStudy (anonymous):

yep, so my answer is c:)

OpenStudy (cwrw238):

right

OpenStudy (anonymous):

thank you

OpenStudy (cwrw238):

yw